jigglypuff314 (jigglypuff314):

and so I was trying to do some practice trig anti derivatives and now my head's spinning >,< a little help pwease? :3 \[\large \int\limits \frac{ dx }{ 1 + e^x } \]

jigglypuff314 (jigglypuff314):

Here's what I've tried doing do far \[\rightarrow \int\limits \frac{(e^{x/2}-e^{x/2}+1)}{1+e^x}dx~\rightarrow \int\limits \frac{e^{x/2}}{1+e^x}dx- \int\limits \frac{e^{x/2}}{1+e^x}dx+ \int\limits \frac{1}{1+e^x}dx\]

Parth (parthkohli):

I haven't studied integration that deeply, but from what you've done so far, you're ending up with the same integral in the end.

jigglypuff314 (jigglypuff314):

I was just trying to find a way to use a trig integral :/ cuz the \[\int\limits \frac{e^{x/2}}{1+e^x}dx \rightarrow \int\limits \frac{e^{x/2}}{(1)^2+(e^{x/2})^2}dx\]

Parth (parthkohli):

Yeah, I see that, but we still don't have a way to compute the last term. But well - both of us know that and I haven't added anything significant so far. Heh.

jigglypuff314 (jigglypuff314):

which is why I'm stuck >,<

OpenStudy (turingtest):

tan(u)=e^(x/2) ?

jigglypuff314 (jigglypuff314):

mmm? I yeah I was thinking u = e^(x/2)

OpenStudy (turingtest):

oh i see that is not where you are stuck

OpenStudy (turingtest):

parth is right that the last term will mess us up

jigglypuff314 (jigglypuff314):

so to try another way entirely?

OpenStudy (turingtest):

yeah

OpenStudy (turingtest):

im super rusty lol, I have to remember how to do this

jigglypuff314 (jigglypuff314):

my math teacher hinted about adding and subtracting something to the numerator to make it work but the wolframalpha solution doesn't do that and it confusing me >,<

OpenStudy (turingtest):

add and subtract e^x in the numerator no trig needed

Parth (parthkohli):

Ooo, how did I not think of that?

jigglypuff314 (jigglypuff314):

\[\int\limits \frac{e^x-e^x+1}{1+e^x}dx\]? then...?

jigglypuff314 (jigglypuff314):

oh if u = e^x then du = e^x dx

Parth (parthkohli):

\(e^x\) and \(1-e^x\) are to be separated.

Parth (parthkohli):

yeah, then a sub.

jigglypuff314 (jigglypuff314):

what do I do with the\[\int\limits \frac{1-e^x}{1+e^x}\]

OpenStudy (turingtest):

other way fellas 1+e^x and -e^x need to be separated

jigglypuff314 (jigglypuff314):

*facedesk* I see now xD

Parth (parthkohli):

Oh, crap. The effects of sleep-deprivation are showing up early (or late?)

OpenStudy (turingtest):

It happens to the best of us... and also to me :P

jigglypuff314 (jigglypuff314):

Thanks for your help! @TuringTest @ParthKohli ^_^

Parth (parthkohli):

Alright, bring 'em on.

Parth (parthkohli):

The integrals, I mean. Can't play the sidekick anymore. :|